A light bulb connected to a 9-V battery has a current of 1.5 A flowing. Find...

  1. A light bulb connected to a 9-V battery has a current of 1.5 A flowing.
    1. Find the resistance of the light bulb (in ohms)
    2. Find the power drawn by the light bulb (in W)
Answer #1

The ohms law states that The potential difference across an ideal conductor is proportional to the current through it.

Ohm's formula is I= V/R

Where I = current in amps, V = voltage in volts, and R = resistance in ohms

When V =9 and I =1.5

then 1.5= 9/R

R= 9/1.5=90/15

R =6 ohms

Power is defined as the rate at which electrical energy is supplied to a circuit  

The formula for Power is  

P = V X I

When V= 9 and I =1.5

P =9 X1.5

=13.5 W